# Hawking radiation in a charged, cosmological black hole

### v1

@@ VI Conclusions
There are four main extremities in the RNdS parameter space: Schwarzschild-de Sitter, Nariai, the lukewarm family and the extremum of the cold regime. An observer of positive energy was released from the maximum of the horizon function, either infalling towards $r_{+}$ or outfalling towards $r_{c}$. The outfaller always sees positive effective temperatures, eventually finding an asymptotic state of thermal equilibrium in pure de Sitter space. The infaller generally perceives positive effective temperatures until they reach a point inbetween $r_{+}$ and $r_{-}$ where the Cauchy horizon induces a blueshift, causing the observer to perceive a negative effective temperature. An exception is at the extremal cold limit, where the observer never sees a negative effective temperature in the ingoing modes. For values of the charge that are large enough, it is possible for the observer to see a blueshift in the outoing modes before they reach the event horizon. The outgoing modes are always negative and generally diverge to infinity at the inner horizon.
To reinforce the generality of the divergence, a density plot of $\log(\kappa^{+}_{i}\Delta)$ evaluated at the inner horizon was plotted over the entire permissible parameter range of the RNdS metric. A finite value of $\kappa^{+}_{i}(r_{-})$ causes the product $\kappa^{+}_{i}\Delta$ to go to zero. The product was finite and non zero over the interior of the sharkfin indicating that the effective temperatures were negatively divergent. However, in accordance with semi-classical predictions, the effective temperature at the inner horizon was seen to be finite for cold and ultracold black holes. Differing from the predictions of [42], the effective temperature still diverges to infinity in the lukewarm regimes. The effective temperature formalism is seen to be one step more accurate in locating regions of possible violation than an examination of the quasinormal modes at the inner horizon. Yet, it is still a semi-classical formulation that does not have the robustness of the quantum stress-energy tensor approach, which further narrows down the violation to the ultracold regime.
@@
In contrast to a cosmological expansion is the case of a cosmological contraction, with a black hole set in an Anti-de Sitter space. Unlike for dS spacetime, AdS features no cosmological horizon, with its outer boundary being infinity. Outgoing modes are sourced from the past horizon, reflect off the boundary at infinity and travel back inwards as ingoing modes. An observer, starting from an arbitrarily chosen rest point will see the outgoing modes get blueshifted under the effect of the contraction, with the mass of the black hole taking predominance, slowly causing the modes to redshift. The outgoing modes remain redshifted until they once again undergo a blueshift effect under the influence of the Cauchy horizon. Yet again, there is an infinite divergence at the inner horizon.
Finally, the validity of the adiabatic approximation was tested by plotting the adiabatic control function against the position of the observer. It is seen, that in general, adiabaticity does not hold, except around a select set of regions. An observer in regimes where adiabaticity is not fulfilled will still see radiation—but it will not be a thermal black body spectrum. While the function $\kappa(r)$, is not affected, the physical conceptualization that accompanies it is no longer that of a “temperature” in conventional sense. But the observer will still see nonzero particle production. A more rigorous calculation of the particle spectrum observed, in regimes where the spectrum is non-thermal, may be accomplished by calculating the scattering coefficients of the modes as seen by an observer in the vacuum state of the emitter [58]. This calculation, while not covered in this paper, is a worthwhile attempt reserved for future work.

### v2

@@ VI Conclusions
There are four main extremities in the RNdS parameter space: Schwarzschild-de Sitter, Nariai, the lukewarm family and the extremum of the cold regime. An observer of positive energy was released from the maximum of the horizon function, either infalling toward $r_{+}$ or outfalling toward $r_{c}$. The outfaller always sees positive effective temperatures, eventually finding an asymptotic state of thermal equilibrium in pure de Sitter space. The infaller generally perceives positive effective temperatures until they reach a point in between $r_{+}$ and $r_{-}$ where the Cauchy horizon induces a blueshift, causing the observer to perceive a negative effective temperature. An exception is at the extremal cold limit, where the observer never sees a negative effective temperature in the ingoing modes. For values of the charge that are large enough, it is possible for the observer to see a blueshift in the outoing modes before they reach the event horizon. The outgoing modes are always negative and generally diverge to infinity at the inner horizon.
To reinforce the generality of the divergence, a density plot of $\log(\kappa^{+}_{i}\Delta)$ evaluated at the inner horizon was plotted over the entire permissible parameter range of the RNdS metric. A finite value of $\kappa^{+}_{i}(r_{-})$ causes the product $\kappa^{+}_{i}\Delta$ to go to zero. The product was finite and nonzero over the interior of the sharkfin indicating that the effective temperatures were negatively divergent. However, in accordance with semiclassical predictions, the effective temperature at the inner horizon was seen to be finite for cold and ultracold black holes. Differing from the predictions of [45], the effective temperature still diverges to infinity in the lukewarm regimes. The effective temperature formalism is seen to be one step more accurate in locating regions of possible violation than an examination of the quasinormal modes at the inner horizon. Yet, it is still a semiclassical formulation that does not have the robustness of the quantum stress-energy tensor approach, which further narrows down the violation to the ultracold regime.
@@
In contrast to a cosmological expansion is the case of a cosmological contraction, with a black hole set in an anti-de Sitter space. Unlike for dS spacetime, AdS features no cosmological horizon, with its outer boundary being infinity. Outgoing modes are sourced from the past horizon, reflect off the boundary at infinity and travel back inward as ingoing modes. An observer, starting from an arbitrarily chosen rest point will see the outgoing modes get blueshifted under the effect of the contraction, with the mass of the black hole taking predominance, slowly causing the modes to redshift. The outgoing modes remain redshifted until they once again undergo a blueshift effect under the influence of the Cauchy horizon. Yet again, there is an infinite divergence at the inner horizon.
Finally, the validity of the adiabatic approximation was tested by plotting the adiabatic control function against the position of the observer. It is seen, that in general, adiabaticity does not hold, except around a select set of regions. An observer in regimes where adiabaticity is not fulfilled will still see radiation—but it will not be a thermal black body spectrum. While the function $\kappa(r)$, is not affected, the physical conceptualization that accompanies it is no longer that of a “temperature” in conventional sense. But the observer will still see nonzero particle production. A more rigorous calculation of the particle spectrum observed, in regimes where the spectrum is nonthermal, may be accomplished by calculating the scattering coefficients of the modes as seen by an observer in the vacuum state of the emitter [63]. This calculation, while not covered in this paper, is a worthwhile attempt reserved for future work.